# Image as an IMU: Estimating Camera Motion from a Single Motion-Blurred Image

### v3

@@ Abstract
In many robotics and VR/AR applications, fast camera motions cause a high level of motion blur, causing existing camera pose estimation methods to fail. In this work, we propose a novel framework that leverages motion blur as a rich cue for motion estimation rather than treating it as an unwanted artifact. Our approach works by predicting a dense motion flow field and a monocular depth map directly from a single motion-blurred image. We then recover the instantaneous camera velocity by solving a linear least squares problem under the small motion assumption. In essence, our method produces an IMU-like measurement that robustly captures fast and aggressive camera movements. To train our model, we construct a large-scale dataset with realistic synthetic motion blur derived from ScanNet++v2 and further refine our model by training end-to-end on real data using our fully differentiable pipeline. Extensive evaluations on real-world benchmarks demonstrate that our method achieves state-of-the-art angular and translational velocity estimates, outperforming current methods like MASt3R and COLMAP.
@@ 1 Introduction
Given a sequence of images, visual odometry (VO) and Structure from Motion (SfM) methods have advanced to the point where they can accurately estimate camera poses even under moderately challenging conditions. However, these methods assume that the camera remains largely stationary during exposure, allowing each image to be treated as a snapshot of the scene that can be matched against subsequent images to compute relative camera poses. This assumption, however, ignores the continuous nature of camera motion and becomes especially problematic during fast movements where motion blur can severely degrade accuracy. Conventional approaches either discard motion-blurred frames or use inertial measurement units (IMUs) to improve robustness against blur. However, using IMUs introduces additional challenges such as sensor synchronization and drift. In this work, we therefore take a fundamentally different approach: rather than treating motion blur as an unwanted artifact, we exploit it as a rich source of information about the camera’s motion (see Figure 1).
Our key observation is that the extent and direction of motion blur provide direct cues about the camera’s motion during an exposure. Building on this idea, we propose a model that, using just the motion-blur in an image, can estimate the relative motion that the camera undergoes during the exposure. Our model works in two stages: first we predict the motion flow field and a monocular depth map, then we recover the relative pose of the camera by solving a linear least squares system. Notably, our method does not require any explicit deblurring process, and with a known exposure time, it yields an instantaneous rotational and translation velocity estimate that can be interpreted as an IMU-like measurement.
Since no existing dataset includes all the necessary data to train such a model, we construct our own dataset using a subset of ScanNet++v2 [58] and show how we obtain realistic motion blur that allows for generalization to in-the-wild images. Furthermore, the fully differentiable nature of our method enables us to refine our model using real-world motion-blurred images, even when ground-truth flows and depths are not available for training. We evaluate our model with angular and translational velocity estimates on real-world motion blurred images to test the generalization capabilities of our model. Our method runs in real-time at 30 FPS and obtains significantly more accurate velocity estimates compared to the existing state of the art methods.
@@
A new approach for estimating camera motion that uses the blur present in a frame to provide extremely robust and accurate relative pose estimates
@@
A pipeline for synthesizing the data needed to train the model (i.e. blurry images with ground-truth motion) from standard SLAM datasets
@@
An evaluation of our method on real-world sequences, showing our method obtains state of the art results and is robust to aggressive camera motion
The rest of the paper is as follows. We first discuss relevant works to our method in Section 2. We then introduce preliminary information and provide a high-level intuition of our method in Section 3. In Section 4, we go into technical details about our method. Section 5 shows the dataset construction process. We evaluate our results against baseline methods in Section 6.
@@ 2 Related Works
Motion from Blur. This work is related to motion from blur methods that attempt to extract pixel-wise flow from a single blurred image. Early approaches assumed spatially invariant blur, with [7] using a linear motion model to extract the flow and [35] using a frequency domain approach. Later methods extended these where [10] first extracts an alpha channel matte and uses the edge gradients to estimate object motion, [41] builds on [35] to use the color channels directly while [14] proposed a variational technique and [1] a learned approach. Our focus, however, is not on estimating the flow field but rather the actual camera motion.
Several works aim to recover either rotational or translational camera motion directly from a blurred image. A seminal work in this direction is [21] which proposes to estimate the camera rotation directly from a single motion-blurred image by searching for edgels to estimate the rotation axis, which is used to help compute the blur length and subsequently the rotation. [24] [3] also estimate the camera rotation with assuming negligible translations, while [2] [8] estimate exclusively linear motion. In contrast, we estimate the full 6DoF velocity. Closer to our method, [62] proposes a neural network to estimate the full relative pose from a single motion-blurred image but directly regresses the pose parameters and requires a clean depth map as input.
Our method estimates depths to retrieve the metric-scaled motion. Similarly, [22], [27], and [34] recover 3D geometry in motion-blurred scenes. [27] is the most relevant, being a visual odometry method which computes the homography during exposure time between a sharp and potentially blurry frame. However, this depend on two images and an off-the-shelf deblurring method, whereas we do not require any deblurring for a single input image.
Deblurring. An adjacent line of work is deblurring, which aims to recover a sharp image given a blurred one. [17] [33] [32] use classical energy minimization frameworks to obtain depth and camera pose estimates, similarly to our method, in addition to the deblurred image. However, compared to deep-learned deblurring methods [60], classical methods are frequently slower during test time and underperform in more general conditions. Therefore, we do not benchmark against such methods given the restrictive assumptions made and significant runtime slowdown. Several deep deblurring methods [45] [15] [61] share parallels to our work in also estimating a motion flow field. [56] reformulates removing the rolling-shutter effects as a deblurring problem with rolling shutter global reset. While our motion field estimates could in theory compute a deblur kernel, our goal is to exploit motion cues from blur artifacts rather than remove them.
@@
COLMAP [40] [39] is an incremental-SfM method that is commonly the default choice for estimating camera parameters and scene geometry, while ORB-SLAM [5] and DSO [12] are classic real-time alternatives. While highly accurate in sequences with large overlap, these traditional systems succumb to failures under conditions including few views and motion blur. These failures motivate deep-learning based approaches which are more robust in these difficulties. Several fully-differentiable SfM pipelines have been proposed for greatly improving 3D reconstruction and pose estimation accuracy and robustness [52] [43] [4]. In particular, DUSt3R [54] and MASt3R [23] have shown that directly regressing pointmaps enables incredible performance, which has inspired a line of “-3R”-related works [59] [51] [42] [57] [53]. SLAM and VO methods have also underwent its own deep revolution [9] [55] [47] [44] [37] [31], with DROID-SLAM [47] obtaining accurate SOTA pose estimates by using RAFT [46] for estimating optical flow. Since no prior methods estimate camera motion from a single image, we decide to benchmark against COLMAP, MASt3R, and DROID-SLAM, approximating velocities via finite differences.
@@ 3 Preliminaries
Before introducing our method, we first revisit the image formation process. During the opening of the shutter, photons come into contact with the camera sensor and are collected to capture each pixel’s intensity. If the camera has moved during the exposure time, each sensor pixel receives photons from different parts of the scene and motion-blur is created. The length of the motion blur traces depends on the speed of the camera as well as the exposure time i.e. when the exposure time is lengthened, e.g. in low-illuminated settings, or if the camera moves faster relative to the scene, the blur traces are longer. More formally, let represent a blurry image, such that:
where represents the response of the sensor to incoming photons at timestep , with the exposure time length denoted by , and representing the conversion from linear space to sRGB [36]. In practice, we approximate the creation of a blurry image in a discretized fashion:
where we assume that are the virtual images that existed in exposure time.
The intuition for our method is that each motion-blurred image can be seen as a composite of several “virtual images,” each capturing the camera at a different pose during the exposure. Each individual virtual image contributes a part of the overall blur, and the blur traces therefore act as cues that represent the motion the camera has undergone during the exposure. In essence, these blur traces can be seen as providing virtual correspondences between the first and last virtual image.
We hypothesize that we can extract these “virtual” correspondences, revealing how each point in the scene moved across the image plane and capturing information similar to an optical flow field. Since we assume the scene is rigid, each virtual correspondence can be interpreted as a match between two distinct virtual images representing the camera at the beginning and end of the exposure period (see Figure 1).
Furthermore, by estimating a depth map from the blurred image, we can combine the pixel motion information extracted from the motion blur traces with the scene geometry to solve for the relative pose change from the start to the end of the exposure.
@@ 4 Method
An overview of our method can be seen in Figure 2. Our approach consists of two stages, the first stage takes as input a single frame and predicts a flow-field representing the motion the camera underwent during the exposure, and a monocular depth map. The second stage uses the depth and flow to solve for the instantaneous velocity of the camera using a differentiable least squares solver.
@@ 4.1 Flow and Depth Prediction
We follow GeoCalib [50] and use SegNeXt [16] as the backbone for our network. The blurry image is passed through a shared SegNeXt encoder that maps to separate decoders, outputting the pixel-wise flow field and depth map . Each flow vector is defined to be the pixel displacement from a given pixel location in the original virtual frame :
such that , and are the corresponding pixel coordinates from the first and last virtual frames and respectively.
We impose an L1 loss for both the flow field and the depth predictions:
where represents the ground truth labels, are weights to balance the two losses, and is the reorientation function defined as:
@@ 4.2 Differentiable Velocity Computation
We can now use and to estimate the relative translation and rotation parameters and across exposure. To do so, we adapt the motion field equations described in Trucco and Verri [48] (see Supplementary 1), which we express as:
@@
This approach leverages the relationship between pixel displacements and camera pose, enabling closed-form estimation of the relative pose parameters and from the predicted flow and depths. With a provided exposure time, we can then compute the instantaneous velocities and across exposure.
@@ 4.3 Direction Disambiguation
While the method we have proposed so far can estimate the magnitude of the camera velocity directly, the direction has a ambiguity because the blurring process erases temporal cues. For instance, both leftward or rightward motion of the camera will result in the same horizontal blur traces. To address this issue, we introduce a simple strategy to resolve the directional ambiguity. Recall is the exposure time length and let be the time gap between the start of frame shutter. For frame , we linearly extrapolate the flow field and apply the warping function to :
For both the forward and backwards flow fields and , we compute the photometric error between the corresponding warped image with the next true frame in the video sequence :
@@
This straightforward photometric error check enables us to reorient the direction of the camera velocities in the correct orientation.
@@ 5 Dataset Curation
| MASt3R [54] [23] | 5.30 / 2.85 / 4.45 | 3.70 / 3.75 / 3.26 | 2.36 / 0.84 / 1.67 | 4.78 / 3.03 / 6.21 | 4.04 / 2.62 / 3.90 | |
| DROID-SLAM [47] | 5.39 / 3.33 / 5.31 | 3.01 / 5.89 / 3.57 | 2.92 / 1.20 / 1.98 | 6.33 / 4.90 / 5.56 | 4.41 / 3.83 / 4.10 | |
@@
| MASt3R [54] [23] | 2.59 / 1.50 / 3.52 | 1.28 / 1.94 / 2.49 | 0.82 / 0.55 / 1.22 | 1.72 / 2.16 / 1.46 | 1.60 / 1.54 / 2.17 | |
| DROID-SLAM [47] | 3.28 / 2.00 / 2.65 | 2.23 / 1.47 / 1.25 | 1.44 / 0.65 / 0.91 | 1.98 / 2.40 / 1.23 | 2.23 / 1.63 / 1.51 | |
@@ 5.1 Synthetic Dataset
To train our model, we require a dataset of blurred images with corresponding ground-truth flows, depths, and relative poses across exposure time. Since no public dataset meets these requirements, we create our own by adapting a subset of ScanNet++v2 [58]. Figure 3 shows our data pipeline.
The ScanNet++v2 dataset consists of about 1000 iPhone videos of various indoor scenes with COLMAP [39] [40] and ARKit-captured poses. Because we are interested in obtaining relative poses that are across very short time spans (fractions of a second) and robust to fast-motion, we opt to use the ARKit poses to compute our ground truth motion. Before curating the dataset, we first preprocess the ARKit poses and filter out any “jumps” in the localization.
@@
We now detail how we select our virtual frames for creating a motion blurred image . For a selected RGB image in a given video sequence, we obtain the following images to have a collection of images . We then use an off-the-shelf frame interpolation model RIFE [18] to interpolate between each consecutive frame, resulting in a total of virtual frames . We follow Equation 2 and first convert in linear space before averaging all images and converting back to sRGB space to create . Rather than simply interpolating virtual frames between two images, we intentionally interpolate between several real-world frames to 1) have more realistic motion trajectories to facilitate generalization for in-the-wild motion blur, and 2) to prevent RIFE from producing unreasonable blurred images. We curate images with for small motion and for large motion.
@@
We only use the first and last virtual frames and for obtaining the ground-truth flow field. We use the low-resolution ARKit depth and in PromptDA [25] to obtain our dense, high-resolution depth map as ground truth. For all pixels, we then perform which uses to backproject all pixel points in and subsequently project them into . Afterwards, we retrieve the flow field with the displacements of the projected points as expressed in Equation 3.
In total, we synthesize about 120k training samples and 1.2k validation samples across 150 ScanNet++v2 sequences, with corresponding blurred images, depth maps and flow fields.

### v4

@@ Abstract
In many robotics and VR/AR applications, fast camera motions lead to a high level of motion blur, causing existing camera pose estimation methods to fail. In this work, we propose a novel framework that leverages motion blur as a rich cue for motion estimation rather than treating it as an unwanted artifact. Our approach works by predicting a dense motion flow field and a monocular depth map directly from a single motion-blurred image. We then recover the instantaneous camera velocity by solving a linear least squares problem under the small motion assumption. In essence, our method produces an IMU-like measurement that robustly captures fast and aggressive camera movements. To train our model, we construct a large-scale dataset with realistic synthetic motion blur derived from ScanNet++v2 and further refine our model by training end-to-end on real data using our fully differentiable pipeline. Extensive evaluations on real-world benchmarks demonstrate that our method achieves state-of-the-art angular and translational velocity estimates, outperforming current methods like MASt3R and COLMAP.
@@ 1 Introduction
Given a sequence of images, visual odometry (VO) and Structure from Motion (SfM) methods can accurately estimate camera poses even in moderately challenging conditions. However, these methods assume that the camera remains largely stationary during exposure, treating each image as a snapshot of the scene that can be matched against subsequent images to compute relative camera poses. This assumption, however, ignores the continuous nature of camera motion and becomes especially problematic during fast movements where motion blur can severely degrade accuracy. Conventional approaches either discard motion-blurred frames or use inertial measurement units (IMUs) to improve robustness against blur. However, using IMUs introduces additional challenges such as sensor synchronization and drift. In this work, we therefore take a fundamentally different approach: rather than treating motion blur as an unwanted artifact, we exploit it as a rich source of information about the camera’s motion (see Figure 1).
Our key insight is that the extent and direction of motion blur provide direct cues about the camera’s motion. Building on this, we propose a model that, using just the motion-blur in an image, can estimate the relative motion that the camera undergoes during the exposure period. Our model works in two stages: first we predict the motion flow field and a monocular depth map, then we recover the relative pose of the camera by solving a linear least squares system. Notably, our method does not require any explicit deblurring process, and with a known exposure time, it yields an instantaneous rotational and translational velocity estimate that can be interpreted as an IMU-like measurement.
Since no existing dataset includes all the necessary data to train such a model, we construct our own dataset using a subset of ScanNet++v2 [yeshwanthliu2023scannetpp] and show how we obtain realistic motion blur that generalizes to in-the-wild images. Furthermore, the fully differentiable nature of our method enables us to refine our predictions using real motion-blurred images, even when flows and depths are unavailable for training. We evaluate our model’s angular and translational velocity estimates on real-world motion blurred images to test the generalization of our model. Our method runs in real-time and obtains more accurate velocity estimates compared to the existing state-of-the-art (SOTA) methods.
@@
A new approach for estimating camera motion that uses the blur present in a frame to provide robust and accurate velocity estimates
@@
A pipeline for synthesizing training data (blurry images with ground-truth motion) from standard vision datasets
@@
An evaluation of our method on real-world sequences, showing our method’s SOTA results and robustness to aggressive camera motion
The rest of the paper is as follows. We first discuss relevant works to our method in Section 2. We then introduce preliminaries and give high-level intuition of our method in Section 3. In Section 4, we describe our method in more detail. Section 5 shows the dataset construction process. We evaluate our results against baseline methods in Section 6.
@@ 2 Related Works
Motion from Blur. This work is related to motion from blur methods that attempt to extract pixel-wise flow from a single blurred image. Early approaches assumed spatially invariant blur, with [chen_motionsmear_1996] using a linear motion model and [rekleitis_motion_1996] using a frequency domain approach. Later methods extended these where [dai_motion_2008] extracts an alpha channel matte and uses the edge gradients to estimate object motion, [schoueri_optical_2009] builds on [rekleitis_motion_1996] to use the color channels directly while [gast_parametric_2016] proposed a variational technique and [argaw_optical_2021] a learned approach. Our focus, however, is not on estimating the flow field but rather the actual camera motion.
Several works aim to recover either rotational or translational camera motion directly from a blurred image. A seminal work in this direction is [Klein2005ASV] which proposes estimating the camera rotation directly from a single motion-blurred image by searching for edgels to estimate the rotation axis, which is used to compute the blur length and subsequently the rotation. [li2006rot-estimate] [Boracchi2008EstimatingCR] also estimate the camera rotation with assuming negligible translations, while [Boracchi2009dir3d] [cortes-osorio_velocity_2019] estimate exclusively linear motion. In contrast, we estimate the full 6DoF velocity. Closer to our method, [zhao2023_motion-from-blur] proposes a neural network to estimate the full relative pose from a single motion-blurred image but directly regresses the pose parameters and requires a clean depth map as input.
Our method estimates depths to retrieve the metric-scaled motion. Similarly, [lee_dense_2013], [liu_mba-vo_2021], and [qiu2019cvpr_worldfromblur] recover 3D geometry in motion-blurred scenes. [liu_mba-vo_2021] is the most relevant, being a visual odometry method which computes the homography during exposure time between a sharp and potentially blurry frame. However, this depends on two images and an off-the-shelf deblurring method, whereas we do not require any deblurring for a single input image.
Deblurring. An adjacent line of work is deblurring, which recovers a sharp image given a blurred one. Similar to our method, [hu2014_jointdepth] [Park2017JointEO] [pan2019_single] also estimate depths and camera poses but use classical energy minimization frameworks to compute them along with the deblurred image. Compared to deep-learned deblurring methods [Zhang2022DeepID], classical methods are notably slower during test time and underperform in more general conditions. Thus, we do not benchmark against such methods given their restrictive assumptions and large runtime slowdown. Several learned deblurring methods [Sun2015LearningAC] [gong2017blur2mf] [zhang_exposure_2022] share parallels to our work in estimating a motion flow field. Related, [Wang_2022_CVPR] reformulates removing rolling-shutter effects as a deblurring problem with rolling shutter global reset. While our motion field estimates could in theory compute a deblur kernel, we aim to exploit motion cues from blur rather than remove it.
@@
COLMAP [schoenberger2016mvs] [schoenberger2016sfm] is commonly the default SfM choice for estimating camera parameters and scene geometry, while ORB-SLAM [orbslam3] and DSO [Engel-et-al-pami2018] are classic real-time alternatives. While highly accurate in sequences with large overlap, these systems succumb to failures under conditions including sparse views and motion blur. These failures motivate learning based approaches which are more robust in these scenarios. Several fully-differentiable SfM pipelines have been proposed for greatly improving 3D reconstruction and pose estimation accuracy and robustness [wang2023vggsfm] [smith24flowmap] [brachmann2024acezero]. In particular, DUSt3R [dust3r_cvpr24] and MASt3R [mast3r_arxiv24] shown how scaling pointmap regression yields SOTA performance in challenging sparse-overlap settings, inspiring a line of “-3R”-related works [zhang2024monst3r] [wang2024spann3r] [smart2024splatt3r] [Yang_2025_Fast3R] [cut3r]. SLAM and VO methods have also underwent its own deep learning revolution [Czarnowski2020deepfactors] [tartanvo2020corl] [teed2021droid] [Sucar_2021_imap] [Rosinol20icra-Kimera] [murai2024mast3rslamrealtimedenseslam], with DROID-SLAM [teed2021droid] obtaining SOTA pose estimates by using RAFT [teed_raft_2020] for estimating optical flow. Since no prior methods estimate camera motion from a single image, we decide to benchmark against COLMAP, MASt3R, and DROID-SLAM, approximating velocities via finite differences.
@@ 3 Preliminaries
Before introducing our method, we first revisit the image formation process. During the opening of the shutter, photons come into contact with the camera sensor and are collected to capture each pixel’s intensity. If the camera has moved during the exposure time, each sensor pixel receives photons from different parts of the scene and motion-blur is created. The length of the motion blur traces depends on the speed of the camera as well as the exposure time; that is, when the exposure time is lengthened (e.g. in low-light settings) or when the camera moves faster relative to the scene, the blur traces become longer. More formally, let represent a blurry image, such that:
where represents the response of the sensor to incoming photons at timestep , the exposure time length denoted by , and representing the linear space to sRGB conversion [rim_2022_ECCV]. In practice, we approximate the blurring process in a discretized fashion:
such that are the “virtual images” during the exposure time. In other words, the true image is the average of all linear space virtual images within exposure time.
The intuition for our method is that each motion-blurred image can be seen as a composite of several virtual images, each capturing the camera at a different pose during the exposure. Each individual virtual image contributes a part of the overall blur, and the blur traces therefore act as cues that represent the motion the camera has undergone during the exposure. In essence, these blur traces can be seen as providing “virtual correspondences” between the first and last virtual image. We hypothesize that we can extract these virtual correspondences, revealing how each point in the scene moved across the image plane and capturing information similar to an optical flow field. By assuming a rigid scene, each virtual correspondence can be interpreted as a match between two virtual images representing the camera at the beginning and end of the exposure period (see Figure 1).
Furthermore, by estimating a depth map from the blurred image, we can combine the pixel motion extracted from the motion blur traces with the scene geometry to solve for the relative pose from the start to end of the exposure.
@@ 4 Method
An overview of our method can be seen in Figure 2. Our approach consists of two stages. The first stage takes as input a single frame and predicts a flow field representing the motion the camera underwent during the exposure and a monocular depth map. The second stage uses the depth and flow to solve for the instantaneous velocity of the camera using a differentiable least squares solver.
@@ 4.1 Flow and Depth Prediction
We follow GeoCalib [veicht2024geocalib] and use SegNeXt [guo2022segnext] as the backbone for our network. The blurry image is passed through a shared SegNeXt encoder that maps to separate decoders, outputting the pixel-wise flow field and depth map . Each flow vector is defined to be the pixel displacement from a given pixel location in the original virtual frame :
such that , the normalized pixel coordinates with respect to the center of the image, and are the corresponding pixel coordinates from the first and last virtual frames and respectively.
We impose an L1 loss for the flow and depths:
where represents the ground truth labels, are weights to balance the losses, and is the reorientation function defined as:
@@ 4.2 Differentiable Velocity Computation
We can now use and to estimate the relative translation and rotation parameters and across exposure. To do so, we adapt the motion field equations described in Trucco and Verri [trucco1998_3dcv] (see supplementary for derivation), which we express as:
@@
This approach leverages the relationship between pixel displacements and camera pose, enabling a closed-form estimate of the relative pose parameters and from the flow and depths. With the exposure time given, we can then compute the instantaneous velocities and across exposure.
@@
Note that our current formulation assumes that the relative pose can be described with a single rigid transform, and we do not model for rolling shutter or non-uniform motion within exposure. However, this is still a reasonable assumption as we later show our results robust to real-world nonlinear blur and rolling shutter effects.
@@ 4.3 Direction Disambiguation
While we can estimate the magnitude of the camera velocity directly, the direction has a ambiguity because the blurring process erases temporal cues. For instance, both leftward and rightward motion of the camera will result in the same horizontal blur traces. To address this, we introduce a simple and fast heuristic to resolve the directional ambiguity. Recall is the exposure time length and let be the inter-frame shutter start interval. Since the motion is small between frames, we first approximate the flow field to be linear. Then, for frame , we linearly extrapolate and apply the warping function to :
Because warping with is linear, warping the real frame is effectively warping each virtual frame and averaging them. Therefore, we can use , the flow between the first virtual frames of and , to warp the true frame (Eq. 13). We use both the forward and backwards flow fields and as in Eq. 12 to compute the extrapolated flow fields and , respectively. Both and are used to compute the photometric error between the corresponding warped image with the next true frame in the video sequence :
@@
In summary, we warp the image in the predicted and opposite flow directions, where we compare the photometric error to decide which direction is consistent with the video frames. This heuristic enables us to quickly reorient the direction of the camera velocities in a consistent orientation.
@@ 5 Dataset Curation
| MASt3R [dust3r_cvpr24] [mast3r_arxiv24] | 5.30 / 2.85 / 4.45 | 3.70 / 3.75 / 3.26 | 2.36 / 0.84 / 1.67 | 4.78 / 3.03 / 6.21 | 4.04 / 2.62 / 3.90 | |
| DROID-SLAM [teed2021droid] | 5.39 / 3.33 / 5.31 | 3.01 / 5.89 / 3.57 | 2.92 / 1.20 / 1.98 | 6.33 / 4.90 / 5.56 | 4.41 / 3.83 / 4.10 | |
@@
| MASt3R [dust3r_cvpr24] [mast3r_arxiv24] | 2.59 / 1.50 / 3.52 | 1.28 / 1.94 / 2.49 | 0.82 / 0.55 / 1.22 | 1.72 / 2.16 / 1.46 | 1.60 / 1.54 / 2.17 | |
| DROID-SLAM [teed2021droid] | 3.28 / 2.00 / 2.65 | 2.23 / 1.47 / 1.25 | 1.44 / 0.65 / 0.91 | 1.98 / 2.40 / 1.23 | 2.23 / 1.63 / 1.51 | |
@@ 5.1 Synthetic Dataset
To train our model, we require a dataset of blurred images with corresponding ground-truth flows, depths, and relative poses across exposure. Since no public dataset meets these requirements, we create our own by adapting a subset of ScanNet++v2 [yeshwanthliu2023scannetpp]. Figure 3 shows our data pipeline. ScanNet++v2 consists of about 1000 iPhone videos of indoor scenes with COLMAP [schoenberger2016sfm] [schoenberger2016mvs] and ARKit-captured poses. We are interested in obtaining relative poses across short time spans (about 0.01 s) and robust to fast-motion, so we opt to use filtered ARKit poses as our ground truth.
We scope this paper to the indoor domain to highlight scenes more susceptible to blur and noise. We do not train on outdoor scenes, so we do not expect our performance to be as accurate outdoors as it is indoors. However, our data pipeline is flexible in that it can easily use any video to generate motion-blurred images, therefore allowing us to obtain images in different domains when desired.
@@
We now show how we create a motion blurred image for training. For a selected RGB image in a video, we obtain the following images to have in total images . We then use an off-the-shelf frame interpolation model RIFE [huang_rife_2022] to interpolate between each consecutive frame, resulting in a total of virtual frames . We follow Equation 2 and first convert in linear space before averaging all images and converting back to sRGB space to create . Even at 60 FPS, frames are sampled too slowly to be averaged for rendering realistic motion blur, so we further interpolate between real-world frames to generate more realistic motion trajectories and facilitate better generalization to in-the-wild motion blur. We curate images with for small motion and for large motion. We emphasize that while we average multiple frames to generate synthetic motion-blurred images, during training/inference our model only ever receives one image at a time.
@@
We only use the first and last virtual frames and for obtaining the ground-truth flow field. We use the low-resolution ARKit depth and in PromptDA [lin2024promptda] to obtain our dense, high-resolution depth map as ground truth. For all pixels, we then perform which uses to backproject all pixel points in and subsequently project them into . Afterwards, we retrieve the flow field with the displacements of the projected points as expressed in Equation 3.
In total, we synthesize about 120k training and 1.2k validation samples across 150 ScanNet++v2 sequences, with corresponding blurred images, depth maps and flow fields.